While trying to make the OAF work, I found that the XYCOM switches for MC1/3 have been set in the bad way for awhile. This means that the hardware filters were bypassed and that MC1 & MC3 were moving around too much at high frequency and possibly causing trouble with the locking. I have put them back into the default position.
On Friday, Jenne and I were playing around with turning the dewhitening off/on to see if it efffected the OAF stability. At the time, I didn't pay too much attention to what the state was. Looks like it was in the wrong state (hardware bypassed) when we found it. For the OAF work, we generally want it in that bypassed state, but its bad because it makes noise in the interferometer. The bits in question are bits 16-23 on the XYCOM screen.
I have updated the snapshot and set the screen in the appropriate settings. I used a swept sine measurement to verify the filter state. In the attached plot, green corresponds to XYCOM green and red corresponds to red. |